OpenStudy (anonymous):

Which of the following are the coordinates of the roots of 2x2 - 8 = 0?

OpenStudy (anonymous):

2x^2-8=0

OpenStudy (anonymous):

divide by 2 and get \[x^2-4=0\] \[x^2=4\] \[x=\pm2\]

OpenStudy (anonymous):

Omg. I'm lost. It says coordinates, but I go on wolframalpha and it doesn't give them to me.